The heights of 2-year-old children are normally distributed with a mean of 32 inches and a standard deviation of 1.5 inches. Pediatricians regularly measure the heights of toddlers to determine whether there is a problem. There may be a problem when a child is in the top or bottom 5% of heights. Determine the heights of 2-year-old children that could be a problem.


1
Expert's answer
2022-05-10T06:55:26-0400

Suppose that "X" is a random variable that has a normal distribution with parameters "\mu=32" and "\sigma=1.5". Consider the probability: "P(X\geq\alpha)=\frac{1}{1.5\sqrt{2\pi}}\int_{\alpha}^{+\infty}e^{-\frac12\left(\frac{x-32}{1.5}\right)^2}dx". By substitutions we receive that "P(X\geq\alpha)=0.05" for "\alpha\approx34.47"(it is rounded to two decimal places). Thus, heights that are greater than "\alpha\approx34.47" are problematic. Consider "P(X\leq\beta)=\frac{1}{1.5\sqrt{2\pi}}\int_{-\infty}^{\beta}e^{-\frac12\left(\frac{x-32}{1.5}\right)^2}dx". By substitutions we obtain that heights that are lower than "\beta\approx29.53" (it is rounded to two decimal places) are problematic.

Answer: heights that are lower than "29.53" and higher than "34.47" are problematic. Values are rounded to two decimal places.